# Constitutional Consensus

### v4

@@ 3.2 Protocol Concepts
For a given $P\subset\Pi$, the function leaderP maps integers to $P$. For each first round of a wave it assigns a formal leader $p\in P$, where a first-round $p$-block $b$ is a leader block if $p=\textit{leader}_{P}(\textit{depth}(b))$. We assume the leader function is defined so that it is fair, in that for every wave and every $p\in P$ there is a subsequent wave for which $p$ is the formal leader of its first round.
Finalizing blocks in waves. The consensus protocol commences with a constitutional blocklace and progresses in waves. The protocol strives to finalize a new block in each wave. The first round of a wave includes candidates for finality, the second round includes endorsements of candidates, and the third round includes ratifications.
A first round block is a candidate for finality in two cases—if it is a formal leader block or the only block in its round. These two possibilities typically arise in high- and low-throughput scenarios, respectively. During low throughput, agents quiesce, not producing a first-round block until they have a new transaction; and when an agent produces a first-round block $b$ that does not conflict with any others, all agents produce second and third round blocks that finalize $b$, in which case $b$ is referred to as spontaneous leader. During high throughput, multiple agents may have transactions to send, and the leader function is used to break symmetry, so that only the formal leader sends a leader block in the first round. Finalized leader blocks of both types induce the total ordering of all blocks, as in other DAG-based protocols [17, 18, 23].
A second-round block $b$ of a wave may approve any number of first-round blocks, but endorses (towards finality) at most one of them, defined as follows. Note that the next two definitions are mutually-recursive.
Definition 3.1 (Endorse).
@@
Note that endorsement is an invariant block property, as it depends only on $b^{\prime}]$.
Next, we specify the finality rule. A first-round block $b\in W$ is final if the third round of $W$ has a $\sigma$-supermajority that approves a second-round $\sigma$-supermajority that endorses $b$. Formally:
Definition 3.2 (Ratified, Final, Ordered, Quiescent).
@@
In addition, the $0^{th}$-wave is defined to be quiescent and its constitutional genesis block to be final.
Note that since only first-round blocks are endorsed, they are the only ones that can be ratified or final. Notice our use of the term “finalised” only applies to such leader blocks; the finalisation of leader blocks induces an order on all the blocks they approve, which are then said to be ordered.
We observe that the block properties defined above are monotonic: {observation}[Monotonicity] If a block is ratified or final in some blocklace $B$, it is also ratified or final, respectively, in any blocklace $B^{\prime}\supset B$. Note, however, that quiescence is not monotonic, because conflicting and non-empty blocks may be added to a round.
The next observation asserts that a wave’s final block is unique. {observation}[Uniqueness] A wave can have at most one ratified or final block.
Proof 3.3.
The observation follows from the fact that a second-round block endorses at most one first-round block in any blocklace it is included in, so to have two $\sigma$-supermajorities of the same round endorse two blocks, at least one correct agent must endorse two blocks in this round; to do so this correct agent must equivocate, a contradiction.
Quiescence, low- and high-throughput scenarios. The protocol is initially quiescent. It begins operations and continues to operate as long as agents have payloads to send. If multiple agents send payloads in the same wave, the protocol switches to high-throughput mode. It returns to being quiescent following a quiescent wave.
Round $r+1$ has a single leader block (red), round $r+2$ consists of a supermajority of blocks (blue as an example) pointing only to (and thus endorsing) the first-round leader block, and round $r+3$ consists of a supermajority of blocks (green as an example) each pointing to (and thus approving) all and only the second-round blocks, and thus ratifying the first-round leader block, which is thus finalized by the wave.
Low-throughput good case: The previous wave ending in round $r$ is quiescent, the spontaneous non-empty leader block (red) may be produced at any time after the previous round $r$ has advanced, followed promptly by two rounds of empty blocks, resulting in a finalising and quiescent wave.
High-throughput good case: The formal leader produces a block (red) within $\delta$ after the previous round $r$ is advanced and all (or many) of the wave’s blocks are non-empty.
Figure 1: A good-case wave $w$.
We first consider the good-case scenarios summarized in Figure 1. During low-throughput all waves are quiescent. A wave commences when some agent has a new transaction. This agent issues a (first-round) non-empty block, and all agents follow promptly with two rounds of empty blocks: Second-round blocks that endorse the first-round block, and third-round blocks each approving all second-round blocks, upon which the first-round block becomes final, with the new wave also being quiescent.
During high-throughput, no wave is quiescent, and the good-case scenario unrolls as above except that the first-round block is issued promptly by the wave’s (formal) leader, whether or not it has any payload, and second- and third-round blocks are typically non-empty, with transactions to be ordered by the the next final leader block.
Going beyond the good case, if during high-throughput agents fail to receive a first-round leader block in a timely manner, they issue first-round blocks of their own. Once a super-majority of agents do so, agents can continue to the next round without waiting for a leader block, with blocks by correct agents to be ordered by the leader of some subsequent finalising wave. Similarly, if during low-throughput there are multiple agents issuing first-round blocks, then typically none of the first-round blocks will become final. In this case, the wave is not quiescent and the protocol proceeds as in the high-throughput scenario.
To formalize these progress rules, we define a round $r$ as advanced in a blocklace $B$ if $B$ includes enough blocks to allow round $r+1$ blocks to be added to the blocklace safely:
Definition 3.4 (Advanced Round).
@@
Definition 3.5 (Valid Block).
An ordinary block $b$ is valid if it is well-formed and round $\textit{depth}(b)-1$ is advanced in $[b]$.
Dissemination. The protocol uses two types of dissemination-inducing blocks: nack and inform; other blocks are referred to as ordinary. Dissemination-inducing blocks are not incorporated in the blocklace of their sender or recipient. Hence, they contribute only to the liveness of the protocol and do not affect its safety, which is derived from structural properties of the blocklace.
Unlike Cordial Miners, which incorporates a blocklace-based dissemination protocol with $O(n^{2})$ amortized communication complexity in the good case, Constitutional Consensus achieves for this model $O(n)$ amortized communication complexity in the good case during high-throughput.
Protocols such as TCP achieve the reliable transfer of a sequence of bytes from the sender to the recipient by employing ack and nack messages for transmission control. The purpose of an ack message is to acknowledge the receipt of a datagram, and of a nack message to report an out-of-order datagram to its sender, e.g., in the case of receiving the $(k+1)^{th}$-datagram before receiving the $k^{th}$-datagram.
In the context of the blocklace, an ‘out-of-order’ block is a block received before some block it points to, and the purpose of a nack-block is indeed to report of such an out-of-order block to its creator (which is not necessarily its sender).
@@
The following definition relates to a setting in which an agent $p$ has a blocklace $B$ and it receives a new block $b$. If $b]\not\subseteq B\cup\{b\}$, then $p$ cannot incorporate $b$ into $B$ since some pointers in $b$ are dangling in $B$. These are the tips of $b]\setminus(B\cup\{b\})$. Hence the following definition:
Definition 3.6 (nack-block).
@@
The reason for the asymmetry of having nack-blocks but not ack-blocks is that any ordinary $p$-block functions as an ack-block, as it attests to all the blocks known to $p$ at the time of its creation. nack blocks enable pull-based forwarding, where an agent who knows it misses a block asks for it. In addition, we make use of push-based forwarding to facilitate leader progress using inform-blocks.
Definition 3.7 (inform-Block).
@@
3.3 Single-Epoch Constitutional Consensus Pseudocode
In the Constitutional Consensus protocol, each agent $p$ maintains two sets of blocks: An input buffer $D$ and a local blocklace $B$. An epoch identifier $e$ points to the epoch’s constitutional genesis block. We use $B^{e}$ to denote the blocklace of epoch $e$ in $B$, which is the subset of $B$ that observes $e$. The protocol is event-driven. Algorithm 1 describes a single-epoch Constitutional Consensus, in which the prevailing constitution is fixed throughout the run.
Initialisation. The first epoch is initialised with $D\leftarrow\phi$, $B\leftarrow\{b\}$, $e\leftarrow b$, where $b$ is a constitutional genesis block and $\emph{payload}\leftarrow\bot$. The initialisation of subsequent epochs is discussed in Algorithm 2.
Let $B:=B^{e}$ herein and $r$ be the maximal advanced round of the current $B^{e}$. Receive: Upon receipt of a $q$-block $b$, $q\in P$
@@ 3.3 Single-Epoch Constitutional Consensus Pseudocode
Accept or Nack: For every $q$-block $b\in D$, $q\in P$,
@@
——- then send once a nack-block for $b^{\prime}$ to $q$.
Issue: Issue once a round $r+1$ block if $r+1$ is a second or third round or a first round that follows a wave that is: 1. quiescent and $\emph{payload}\neq\bot$, or 2. non-quiescent and (i) $p=\textit{leader}_{P}(r+1)$ or (ii) $r$ has been advanced for $9\Delta$. Issue Backlog: If $\emph{payload}\neq\bot$ and no round $r$ or round $r+1$ block has been issued by $p$, then issue once a round $r$ block. Inform Leader: If $r$ is a third-round of a non-quiescent wave that has been advanced for $2\Delta$ then send once an inform-block for round $r+1$ to $\textit{leader}_{P}(r+1)$. Output: If $B$ contains a final block $b$ of depth greater than the previous final block then output every block in $\tau(b)$ (Definition 3.8) unless it has already been output.
issue a round $k$ block means create a block $b$ with payload and pointers to the $P$-blocks that are tips of $B_{k-1}$, add $b$ to $B$, send $b$ to every $q\neq p\in P$, and set $\emph{payload}\leftarrow\bot$. judiciously send $[b]$ to $q$ means send every $b^{\prime}\in[b]$ to $q$ unless (i) $p$ has already sent $b^{\prime}$ to $q$, or (ii) $b^{\prime}\in[b_{q}]$ for some $q$-block $b_{q}\in B\cup D$.
@@
Protocol walk-through.
Receive: valid incoming ordinary blocks are inserted to $D$.
Accept or Nack: if any of their predecessors are missing, $p$ waits $\Delta$ to issue a nack, and once a block in $D$ has no missing predecessors it is accepted to $B$ and removed from $D$.
If $r$ is a first or second round in its wave, then $p$ immediately (upon $r$ advancing) issues a round $r+1$ block. If $r$ is a third round, then block issuing (in the ensuing wave) depends on the mode. During low throughput (i.e., following a quiescent wave), $p$ issues a block if and only if it has payload to send. During high-throughput, $p$ issues a block in case (i) it is the formal leader, or (ii) it times-out on the leader. Here, a long timeout is used ($9\Delta$), before which we try to inform the leader that the round has advanced to give it a chance to progress.
Inform Leader: Upon timing out – with a short timeout of $2\Delta$ – on the formal leader in the first round of a high-throughput wave, $p$ sends an inform block to the formal leader in order to cause it to advance in case it is correct.
Output: Once the protocol finalizes a block $b$, it orders the blocks approved by $b$ that have not been ordered yet using the function $\tau$. The definition of $\tau$ is essentially-identical to that of Cordial Miners: It starts from the most-recent final block in $B$ and iterates through ratified blocks. The output of $\tau(b)$ is a sequence that includes all blocks approved by $b$. Formally:
Definition 3.8 ($\tau$).
[…]
When $\tau$ is called with a block $b$, it makes a recursive call with a block $b^{\prime}$ ratified in $[b]$. Note that $b^{\prime}$ is not necessarily final in $[b]$, nor final in the closure of the final block with which $\tau$ was called initially. In the implementation of $\tau$, the results of recursive calls can be cached and used in subsequent calls, saving the need to recompute them.
Proposition 3.9.
@@
In either case, not both $b_{1}$ and $b_{2}$ will be in $\tau(b)$.
Safety and liveness. We assume that the prevailing constitution is safe and live (Definition 2.2). In Section 4 we prove that $\tau$ is monotonic. This ensures that consecutive calls to $\tau$ by a correct agent produce ever-increasing sequences, each a prefix of its successor. Hence, the protocol needs only output the increment of the output sequence of $\tau$ over its output from the earlier call, if any. In addition, as we shall see, during a protocol run the outputs of calls to $\tau$ by correct agents always occur in the same order, which is the basis for protocol safety.
Liveness is achieved after GST. Whenever there is a post-GST wave with a correct leader following a non-quiescent wave, all correct agents accept and endorse the leader’s first-round block (without timing out), and so no correct non-leader agent sends a first-round block. Therefore, no agent can produce a valid second round block that does not endorse the leader block. So the leader block is endorsed by all second round blocks, ratified by all third round blocks, and finalized. Once any agent has payload to send following a quiescent wave, the agent issues a block, and either the block is finalized in the same wave or the wave is non-quiescent and the protocol switches to high-throughput mode.
Latency and complexity. Regarding latency, we note that after GST and in the good case a leader block (formal or spontaneous) is finalized by its wave within $3\delta$, during both low- and high-throughput (See Figure 1).
Regarding communication complexity: During low throughput, in the good case the spontaneous leader block is sent to all, constant-size endorsement blocks are sent all-to-all, and the culprit is the third round, where blocks each with $O(n)$-pointers are sent to all-to-all. In Morpheus [23], this issue is alleviated by having second-round blocks be replaced by threshold signatures on the first block (in low-throughput mode): a $\sigma$-supermajority of these signatures can be amalgamated to form a single signature of constant length, and then third-round blocks can be replaced with threshold signatures on that single signature, giving communication complexity $O(n^{2})$.
A similar idea can be applied here, by having hitherto-empty blocks carry a threshold signature instead. Specifically, a second-round block, instead of being empty, should carry as payload a threshold signature of the pointer to the first-round leader block. A third-round block that observes a supermajority of second-round blocks with threshold signatures, instead of being empty and pointing to all second-round blocks, would carry as payload an amalgamation of these threshold signatures, and point only to its own second-round block. Thus, in the good case, all blocks of a low-throughput wave are of constant size, resulting in $O(n^{2})$ communication complexity. To accommodate this change, the definition of a third-round $p$-block $b$ ratifying a first-round leader block $b^{\prime}$ has to include the case that $b$ points only to the second round $p$-block, but its payload is an amalgamation of threshold signatures of a supermajority of second-round blocks that endorse $b^{\prime}$.
During high-throughput, all blocks carry payloads. If the payloads have at least $O(n)$ transactions each, then, in the good case, the amortized communication complexity per transaction is $O(n)$.
4 Safety and Liveness of Single-Epoch Constitutional Consensus
4.1 Safety
The safety properties of Definition 2.4 can be shown by proving the consistency of the output sequences of correct agents, formally defined as follows: We say that two sequences $x$ and $y$ are consistent if one is a prefix of the other, namely there is a sequence $z$ such that $x\cdot z=y$ or $x=y\cdot z$, with $\cdot$ denoting sequence concatenation.
{observation}
Proof 4.1.
Validity trivially holds because every agent participates in every epoch. Assuming the output sequences of correct agents are all consistent with each other, we take $T$ to be their supremum and get Consistency.
Theorem 4.2.
First, recall the assumption that in a blocklace $B$ created by the protocol with prevailing constitution $(P,\sigma)$, the $P$-blocks in $B$ include equivocations by at most $f=(2\sigma-1)n$ agents. We refer to blocklaces and waves satisfying this property as constitutionally-safe wrt $(P,\sigma)$.
@@ 4.1 Safety
Then the blue supermajority and the finalising $\sigma$-supermajority of $b$, let’s call it the green supermajority, must have a correct agent in their intersection. Namely, there is a correct agent $p\in P$ with a blue $p$-block that must observe a green $p$-block (which may actually be the same block if $W$ and $W^{\prime}$ are consecutive waves), which in turn must observe a $\sigma$-supermajority of blocks that endorse $b$. Since $b^{\prime}$ observes the blue $p$-block that observes the green $p$-block that ratifies $b$, then $b$ is ratified by $b^{\prime}$.
Proposition 4.6 ($\tau$ Monotonicity).
@@
By Observation 3.2, all the blocks final in $B$ are also final in $B^{\prime}$. By Observation 3.2, each wave in $B,B^{\prime}$ includes at most one ratified or final block, hence the ratified and final blocks are totally ordered by their depth. If $B,B^{\prime}$ have the same most-recent final block, or if $B$ has no final block, then the proposition holds trivially. Assume $b\neq b^{\prime}$ are the most-recent final blocks in $B,B^{\prime}$, respectively, so necessarily $b^{\prime}$ is more recent than $b$. Let $b=b_{1},\ldots,b_{k}=b^{\prime}$, $k\geq 2$, be the sequence of all ratified blocks in $B^{\prime}$ between and including $b$ and $b^{\prime}$. By Proposition 4.4 and the definition of $\tau$, $\tau(b^{\prime})$ will eventually call recursively $\tau(b)$ and hence $\tau(b)$ is a prefix of $\tau(b^{\prime})$, implying that $\tau(B)$ is a prefix of $\tau(B^{\prime})$.
{observation}
@@
Proof 4.8.
By monotonicity of $\tau$ over constitutionally-safe blocklaces (Proposition 4.6), $\tau(B)$ and $\tau(B^{\prime})$ are both prefixes of $\tau(B\cup B^{\prime})$, which implies that one is a prefix of the other, namely they are consistent.
Proof 4.9 (Proof of Theorem 4.2).
By assumption, the $P$-blocks in $B\cup B^{\prime}$ include equivocations by at most $f=(2\sigma-1)n$ agents, so this set is constitutionally-safe. Hence, by Observation 4.1, the outputs of any two correct agents running Constitutional Consensus are always consistent. By Observation 4.1, the protocol is safe.
4.2 Liveness

### v5

@@ A.2. Protocol concepts
For a given $P\subset\Pi$, the function leaderP maps integers to $P$. For each first round of a wave it assigns a formal leader $p\in P$, where a first-round $p$-block $b$ is a leader block if $p=\textit{leader}_{P}(\textit{depth}(b))$. We assume the leader function is fair, in that for every wave and every $p\in P$ there is a subsequent wave for which $p$ is the formal leader of its first round.
Definition A.1 (Endorse).
@@
Note that endorsement is an invariant block property, as it depends only on $b^{\prime}]$. Next, we specify the finality rule. A first-round block $b\in W$ is final if the third round of $W$ has a $\sigma$-supermajority that approves a second-round $\sigma$-supermajority that endorses $b$. Formally:
Definition A.2 (Ratified, Final, Ordered, Quiescent).
@@
In addition, the $0^{th}$-wave is quiescent and its constitutional genesis block final.
Note that since only first-round blocks are endorsed, they are the only ones that can be ratified or final. Notice our use of the term “finalised” only applies to such leader blocks; the finalisation of leader blocks induces an order on all the blocks they approve, which are then said to be ordered. We observe that the block properties defined above are monotonic:
Observation A.3 (Monotonicity).
If a block is ratified or final in some blocklace $B$, it is also ratified or final, respectively, in any blocklace $B^{\prime}\supset B$.
To formalize the round progress rules, we define a round $r$ as advanced in a blocklace $B$ if $B$ includes enough blocks to allow round $r+1$ blocks to be added to the blocklace safely:
Definition A.4 (Advanced Round).
@@
Definition A.5 (Valid Block).
A block $b$ is valid if it is well-formed and round $\textit{depth}(b)-1$ is advanced in $[b]$.
Dissemination. The protocol uses two types of dissemination-inducing blocks: nack and inform; other blocks are referred to as ordinary. Dissemination-inducing blocks are not incorporated in the blocklace of their sender or recipient. Hence, they contribute only to the liveness of the protocol and do not affect its safety, which is derived from structural properties of the blocklace. Unlike Cordial Miners, which incorporates a blocklace-based dissemination protocol with $O(n^{2})$ amortized communication complexity in the good case, Constitutional Consensus achieves for this model $O(n)$ amortized communication complexity in the good case during high-throughput.
@@
The following definition relates to a setting in which an agent $p$ has a blocklace $B$ and it receives a new block $b$. If $b]\not\subseteq B\cup\{b\}$, then $p$ cannot incorporate $b$ into $B$ since some pointers in $b$ are dangling in $B$. These are the tips of $b]\setminus(B\cup\{b\})$.
Definition A.6 (nack-block).
@@
nack blocks enable pull-based forwarding, where an agent who knows it misses a block asks for it. In addition, we make use of push-based forwarding to facilitate leader progress using inform-blocks.
Definition A.7 (inform-Block).
@@
A.3. The Ordering Function $\tau$
Definition A.8 ($\tau$).
[…]
When $\tau$ is called with a block $b$, it makes a recursive call with a block $b^{\prime}$ ratified in $[b]$. In the implementation of $\tau$, the results of recursive calls can be cached and used in subsequent calls, saving the need to recompute them.
Let $B:=B^{e}$ herein and $r$ be the maximal advanced round of the current $B^{e}$. Input:: Upon input a transaction $T$, $\emph{payload}\leftarrow\emph{payload}\cup\{T\}$ Receive:: Upon receipt of a $q$-block $b$, $q\in P$
@@ A.3. The Ordering Function $\tau$
Accept or Nack:: For every $q$-block $b\in D$, $q\in P$,
@@
——- then send once a nack-block for $b$ to $q$.
Issue:: Issue once a round $r+1$ block if $r+1$ is a second or third round or a first round that follows a wave that is: (1): quiescent and $\emph{payload}\neq\bot$, or (2): non-quiescent and (i) $p=\textit{leader}_{P}(r+1)$ or (ii) $r$ has been advanced for $9\Delta$. Issue Backlog:: If $\emph{payload}\neq\bot$ and no round $r$ or round $r+1$ block has been issued by $p$, then issue once a round $r$ block. Inform Leader:: If $r$ is a third-round of a non-quiescent wave that has been advanced for $2\Delta$ then send once an inform-block for round $r+1$ to $\textit{leader}_{P}(r+1)$. Output:: If $B$ contains a final block $b$ of depth greater than the previous final block then output all the transactions in every block in $\tau(b)$ (Definition A.8) that have not already been output.
issue a round $k$ block means create a block $b$ with payload and pointers to the tips of $B_{k-1}$, add $b$ to $B$, send $b$ to every $q\neq p\in P$, and set $\emph{payload}\leftarrow\bot$. judiciously send $[b]$ to $q$ means send every $b^{\prime}\in[b]$ to $q$ unless (i) $p$ has already sent $b^{\prime}$ to $q$, or (ii) $b^{\prime}\in[b_{q}]$ for some $q$-block $b_{q}\in B\cup D$.
@@
Appendix B Pseudocode Walkthrough
Receive: valid incoming ordinary blocks are inserted to $D$. Subsequently, the Accept or Nack is applied: if any of their predecessors are missing, $p$ waits $\Delta$ to issue a nack, and once a block in $D$ has no missing predecessors it is accepted to $B$ and removed from $D$.
If $r$ is a first or second round in its wave, then $p$ immediately (upon $r$ advancing) issues a round $r+1$ block. If $r$ is a third round, then block issuing (in the ensuing wave) depends on the mode. During low throughput (i.e., following a quiescent wave), $p$ issues a block if and only if it has payload to send. During high-throughput, $p$ issues a block in case (i) it is the formal leader, or (ii) it times-out on the leader. Here, a long timeout is used ($9\Delta$), before which we inform the leader that the round has advanced through the Inform Leader rule: Upon timing out – with a short timeout of $2\Delta$ – on the formal leader in the first round of a high-throughput wave, $p$ sends an inform block to the formal leader in order to cause it to advance in case it is correct.
Appendix C Latency and Communication Complexity
Latency. After GST and in the good case a leader block (formal or spontaneous) is finalised by its wave within $3\delta$, during both low- and high-throughput (See Figure 1).
Communication complexity. During low throughput, in the good case the spontaneous leader block is sent to all, constant-size endorsement blocks are sent all-to-all, and the culprit is the third round, where blocks each with $O(n)$-pointers are sent to all-to-all. In Morpheus (Lewis-Pye and Shapiro, 2025), this issue is alleviated by having second-round blocks be replaced by threshold signatures on the first block (in low-throughput mode): a $\sigma$-supermajority of these signatures can be amalgamated to form a single signature of constant length, and then third-round blocks can be replaced with threshold signatures on that single signature, giving communication complexity $O(n^{2})$.
A similar idea can be applied here, by having hitherto-empty blocks carry a threshold signature instead. Specifically, a second-round block, instead of being empty, should carry as payload a threshold signature of the pointer to the first-round leader block. A third-round block that observes a supermajority of second-round blocks with threshold signatures, instead of being empty and pointing to all second-round blocks, would carry as payload an amalgamation of these threshold signatures, and point only to its own second-round block. Thus, in the good case, all blocks of a low-throughput wave are of constant size, resulting in $O(n^{2})$ communication complexity. During high-throughput, all blocks carry payloads. If the payloads have at least $O(n)$ transactions each, then, in the good case, the amortized communication complexity per transaction is $O(n)$.
Appendix D Single-Epoch Protocol: Delays
@@
Appendix E Safety and Liveness of Single-Epoch Constitutional Consensus
E.1. Safety
In this section, we prove the single-epoch protocol’s safety.
Theorem E.1.
The safety properties (Consistency and Validity) in Definition 2.3 can be shown by proving the consistency of the output sequences of correct agents, formally defined as follows: We say that two sequences $x$ and $y$ are consistent if one is a prefix of the other, namely there is a sequence $z$ such that $x\cdot z=y$ or $x=y\cdot z$, with $\cdot$ denoting sequence concatenation.
Observation E.2.
Proof.
Validity trivially holds because every agent participates in every epoch. Assuming the output sequences of correct agents are all consistent with each other, we take $T$ to be their supremum and get Consistency. ∎
The next observation asserts that a wave’s final block is unique.
Observation E.3 (Uniqueness).
A wave can have at most one ratified or final block.
Proof.
The observation follows from the fact that a second-round block endorses at most one first-round block in any blocklace it is included in, so to have two $\sigma$-supermajorities of the same round endorse two blocks, at least one correct agent must endorse two blocks in this round; to do so this correct agent must equivocate, a contradiction. ∎
To prove consistency, first, recall the assumption that in a blocklace $B$ created by the protocol with prevailing constitution $(P,\sigma)$, the $P$-blocks in $B$ include equivocations by at most $f=(2\sigma-1)n$ agents. We refer to blocklaces and waves satisfying this property as constitutionally-safe wrt $(P,\sigma)$.
Proposition E.4.
@@ E.1. Safety
In either case, not both $b_{1}$ and $b_{2}$ will be in $\tau(b)$. ∎
@@
Then the blue supermajority and the finalising $\sigma$-supermajority of $b$, let’s call it the green supermajority, must have a correct agent in their intersection. Namely, there is a correct agent $p\in P$ with a blue $p$-block that must observe a green $p$-block (which may actually be the same block if $W$ and $W^{\prime}$ are consecutive waves), which in turn must observe a $\sigma$-supermajority of blocks that endorse $b$. Since $b^{\prime}$ observes the blue $p$-block that observes the green $p$-block that ratifies $b$, then $b$ is ratified by $b^{\prime}$. ∎
Proposition E.7 ($\tau$ Monotonicity).
@@
By Observation A.3, all the blocks final in $B$ are also final in $B^{\prime}$. By Observation E.3, each wave in $B,B^{\prime}$ includes at most one ratified or final block, hence the ratified and final blocks are totally ordered by their depth. If $B,B^{\prime}$ have the same most-recent final block, or if $B$ has no final block, then the proposition holds trivially. Assume $b\neq b^{\prime}$ are the most-recent final blocks in $B,B^{\prime}$, respectively, so necessarily $b^{\prime}$ is more recent than $b$. Let $b=b_{1},\ldots,b_{k}=b^{\prime}$, $k\geq 2$, be the sequence of all ratified blocks in $B^{\prime}$ between and including $b$ and $b^{\prime}$. By Proposition E.6 and the definition of $\tau$, $\tau(b^{\prime})$ will eventually call recursively $\tau(b)$ and hence $\tau(b)$ is a prefix of $\tau(b^{\prime})$, implying that $\tau(B)$ is a prefix of $\tau(B^{\prime})$. ∎
Observation E.8.
@@
Proof.
By monotonicity of $\tau$ over constitutionally-safe blocklaces (Proposition E.7), $\tau(B)$ and $\tau(B^{\prime})$ are both prefixes of $\tau(B\cup B^{\prime})$, which implies that one is a prefix of the other, namely they are consistent. ∎
Proof of Theorem E.1.
By assumption, the $P$-blocks in $B\cup B^{\prime}$ include equivocations by at most $f=(2\sigma-1)n$ agents, so this set is constitutionally-safe. Hence, by Observation E.8, the outputs of any two correct agents running Constitutional Consensus are always consistent. By Observation E.2, the protocol is safe. ∎
E.2. Liveness